OpenStudy (anonymous):

Lea drives from Manayunk to Wilmington at 2/3 of the average speed at which she returned from Wilmington to Manayunk. What percent of her total driving time did it take Lea to drive from Manayunk to Wilmington?

OpenStudy (zale101):

try using the d=rt formula d = distance r = rate t = time

OpenStudy (anonymous):

okay one sec... and that was the formula i couldn't remember... I'll try it out

OpenStudy (zale101):

Okay :)

OpenStudy (anonymous):

well what would be the distance? cause the rate is 2/3?

OpenStudy (anonymous):

or what would be the time?

OpenStudy (anonymous):

Is is one of those tricky question, does if have answer choices? (Just curious).

OpenStudy (anonymous):

*This

OpenStudy (anonymous):

yes it does, hold on, I will post them.

OpenStudy (zale101):

since the total driving time is not giving, it would be difficult to know. Maybe there's a method need to be used to know the time first and use the d=rt formula

OpenStudy (anonymous):

A) 33(1/3)% B) 40% C) 50% D) 60% E) 66(2/3)%

OpenStudy (zale101):

@Mertsj

OpenStudy (anonymous):

hmm idk would time be 1/60 idk, haha

OpenStudy (anonymous):

Hmm, I think Zale is on the right track, you most definitely use \(d=rt\).

OpenStudy (anonymous):

but what would be the time, that is all I need I think... cause I think the rate would be 2/3

OpenStudy (anonymous):

Possible we use direct variation? \(y=kx\) So: \(d=(\dfrac{2}{3})t\) \(k=\dfrac{y}{x}\) \(\dfrac{2}{3}=\dfrac{y}{x}\) I guess we could try the multiple choice from there? I don't know >.<

OpenStudy (anonymous):

no that is not enough information

OpenStudy (anonymous):

I know :-P. I was just attempting it, hopefully Mertsj can explain it.

OpenStudy (mertsj):